OpenStudy (anonymous):

How do I integrate: 1/(1-x^2)^1/2 I tried using u substitution. Is this an integration using partial fractions?

OpenStudy (anonymous):

Put \(x=\sin\theta\)

OpenStudy (anonymous):

ah trig substitution. k will look it up.

OpenStudy (anonymous):

\[dx = d\theta*cos\theta\]Yeah, Trig makes it easy.

OpenStudy (anonymous):

x=sin(@) dx=cos(@)d@ (1/cos@)cos@d@ so @+c arcsinx+c

OpenStudy (anonymous):

Good Job Jamal Ahmed :)

OpenStudy (anonymous):

thank you guys soooo much. Taking a differential equations so I gotta remember alll this stuff again. I'll be back, lol.